OpenStudy (unklerhaukus):

Consider the differential equation: \[y''+\lambda y = 0\] For the given boundary conditions, find the eigenvalues and corresponding eigenfunctions. \[y(0)+y'(0)=0,\qquad y(1)=0\]

OpenStudy (holsteremission):

Assume \(\lambda>0\). The characteristic equation is \(r^2+\lambda=0\), so \(r=\pm i\sqrt\lambda\). Then the characteristic solution is \[y=C_1\cos\sqrt\lambda~t+C_2\sin\sqrt\lambda~t\]and the boundary conditions give the system \[\begin{cases} C_1\cos\sqrt\lambda+C_2\sin\sqrt\lambda=0&\color{lightgray}{y(1)=0}\\[1ex] C_1+C_2\sqrt\lambda=0&\color{lightgray}{y(0)+y'(0)=0} \end{cases}\]Solving both equations for \(\dfrac{C_1}{C_2}\), you get \[\sqrt\lambda=\tan\sqrt\lambda\]so any eigenvalue \(\lambda>0\) satisfies this equation (infinitely many of these). - - - - - - - - - - - - - - - - - - - - - - - - - - - Now assume \(\lambda=0\). The solution would then be \[y=C_1+C_2t\]and the boundary conditions require \[\begin{cases}C_1+C_2=0\\[1ex]C_1+C_2=0\end{cases}\]so there's an infinite number of linear solutions with \(C_1+C_2=0\). - - - - - - - - - - - - - - - - - - - - - - - - - - - Now assume \(\lambda<0\), so the characteristic equation has roots \(r=\pm\sqrt\lambda\) and the solution becomes \[y=C_1\cosh\sqrt\lambda~t+C_2\sinh\sqrt\lambda~t\]with the boundary conditions requiring that \[\begin{cases} C_1\cosh\sqrt\lambda+C_2\sinh\sqrt\lambda=0\\[1ex] C_1+\sqrt\lambda~C_2=0\end{cases}\]Similar to the case where \(\lambda>0\), negative eigenvalues will be the negative roots to \(x-\tanh x\). This only has one root at \(x=0\), so there are no negative eigenvalues satisfying these boundary conditions. - - - - - - - - - - - - - - - - - - - - - - - - - - - I'd have to look into some notes on BVPs to figure out the part about finding the eigenfunctions. Hope this was at least somewhat helpful.

OpenStudy (irishboy123):

i reckon you're crossing into Sturm-Liouville. @Kainui

OpenStudy (unklerhaukus):

\[\ \newcommand\ps[1]{\left(#1\right)}}\] \begin{align*} y''+\lambda y &= 0 &y(0) + y'(0) = 0,\qquad y(1) = 0\\[2ex] m^2+\lambda &= 0 \end{align*} \(\textbf{Case I}\) \(\lambda = 0\), \(m=0,0\) \begin{align*} y &= A+Bx & y(1)&=A+B=0\\ && &B=-A\\ &= A(1-x) & y(0)&=A\\[1ex] y' &= -A & y'(0)&=-A\\[1ex] && y'(0)+y(0)&=A-A=0\\[2ex] y_0 &= 1-x & \lambda_0&=0 \qquad n=0 \end{align*} \(\textbf{Case II}\) \(\lambda = -\mu^2 < 0\), \(m=\pm\mu\) \begin{align*} y &= Ae^{\mu x}+Be^{-\mu x} & y(0) &=A+B\\[1ex] y' &= A\mu e^{\mu x}-B\mu e^{-\mu x} & y'(0)&=(A-B)\mu\\[2ex] && y(0)+y'(0)&=A+B+(A-B)\mu = 0\\ && &A+B = 0\qquad A-B = 0\\ && &A = 0\qquad B = 0\\ y &= 0 \end{align*} \(\textbf{Case III}\) \(\lambda = \mu^2 > 0\), \(m=\pm i\mu\) \begin{align*} y &= A\cos\mu x+B\sin\mu x & y(1) &=A\cos\mu+B\sin\mu=0 \\ && &B= -A\cot\mu\\ &= A\ps{\cos\mu x-\cot\mu\sin\mu x} &y(0) &= A\\[1ex] y' &= -A\mu\ps{\sin\mu x+\cot\mu\cos\mu x} & y'(0)&= -A\mu\cot\mu\\[1ex] && y(0)+y'(0)&= A-A\mu\cot\mu=0\\[2ex] y &= A\ps{\cos\mu_n x-\tfrac1{\mu_n}\sin\mu_n x} &&\mu_n = \tan\mu_n\qquad n = 1,2,3\dots\\[2ex] y_n &= \cos\mu_n x-\tfrac1{\mu_n}\sin\mu_n x &\lambda_n &= \mu_n^2 \end{align*}

OpenStudy (unklerhaukus):

\[\mu_n= 4.493, 7.725,10.904, 14.066, 17.221, \dots \]